87+92+K(all divided by 3)=91 Thanks
Answer by checkley75(3666)

(87+92+K)/3=91
(179+K)/3=91 NOW CROSS MULTIPLY & WE GET:
179+K=273
K=273-179
K=94 ANSWER.